thank you so much all the help, im so grateful!! god bless !
(1 point) Archeologists have discovered a sample of ancient Egyptian skulls from 150 A.C. A sample of 10 Egyptian skull bread 1

Previous Problem Problem List Next Problem (1 point) Archeologists have discovered a sample of ancient Egyptian skulls from 42
Answer #1

x is a v. of skull breadths Exi = 1320 1) 5 = Exi 1320 132 - lo 2 2 s = EX (a)? N-1 برا - N) 174456 (132)? = 24 10-1 10(10-1)